OpenStudy (anonymous):

Iron is produced from its ore by the reaction: 2C+O2 -> 2CO Fe2O3 +3CO -> 2Fe +3CO2 how many moles of O2 are needed to produce 1 mole of Fe?

OpenStudy (anonymous):

3 CO = 2Fe 2C0= 4/3Fe O2 = 4/3 Fe for one mole of Fe we need 3/4 moles of O2

OpenStudy (anonymous):

You can use substitution:\[Fe_2O_3 + 3/2(2C + 0_2) \rightarrow 2Fe + 3CO_2\] And then divide both sides of the equation (by 2 in this case) to end up with 1 mole of iron:\[1/2(Fe_2O_3) + 3/4(2C + 0_2) \rightarrow Fe + 3/2(CO_2)\] You wind up with the same answer shankvee gave:\[3/4(0_2) + other stuff \rightarrow Fe + other stuff\]